a man travels at certain distance at the rate of 70km/hour and return at the same point at the rate of 55 km/hour .What is the average speed of the journey ?
Answer by mananth(16949)

distance = x
rate = 70 km/h
..
time x/70
..
return
time= x/55
...
Total Average rate = total distance/total time
=2x/(x/70+x/55)
=2x/(125x/3850)
=2x*3850/125x
=61.5 km/h
..
There is an easy solution.
let him drive for 1 hour
so forward journey he takes 1 hour.
distance he travels = 70 km
RETURN
distance = 70
rate = 55
time = 70/55= 1.27 hours
Total time taken = 1+1.27= 2.27 hours
total distance = 140 km
rate = 140/2.27
average rate for entire journey = 61.6 km/h
...
m.ananth@hotmail.ca